DIRECT MEASUREMENTS OF THE FUNDAMENTAL ROTATIONAL TRANSITIONS OF CD AND ^{13}CH (X^{2}Π_{r})

D. T. HALFEN, \underline{L. M. ZIURYS}, University of Arizona, Steward Observatory, Department of Chemistry Arizona Radio Observatory, Tucson, AZ 85721; J. C. PEARSON, AND B. J. DROUIN, Jet Propulsion Laboratory, Pasadena, CA 91109.

The N = 1 → 1 and N = 1 → 2 rotational transitions of CD, and the N = 1 → 1 lines of ^{13}CH have been measured in their ^{2}Π_{r} ground electronic states using sub-mm direct absorption spectroscopy. The measurements below 600 GHz (CD: N = 1 → 1 and ^{13}CH: N = 1 → 1) were carried out at Arizona, while those in the 900 GHz range were conducted at JPL (CD: N = 1 → 1). The two radical species were created in an electrical discharge of either ^{13}CH_{4} or CD_{4}. Both lambda-doubling and hyperfine splitttings were resolved in the spectra. The data were analyzed with a case(b) effective Hamiltonian, resulting in an improvement in the lambda-doubling and deuterium, proton, and ^{13}C hyperfine constants. Highly accurate rest frequencies are now available for astronomical searches for these species. CH is an abundant and widespread interstellar molecule, and thus CD and ^{13}CH should be of astrophysical interest.